OpenStudy (anonymous):

y=3/4x-2 Find the slope and y-intercept of the line in the given equation.

OpenStudy (aroub):

the slope=3/ 4 y-intercept=-2

OpenStudy (anonymous):

slope =3/4 y inter=-2

OpenStudy (anonymous):

Okay so how do we get the graph for is

OpenStudy (anonymous):

as in how do u maen?

OpenStudy (anonymous):

i would suggest u sketch it

OpenStudy (anonymous):

He he. I suggest put a dot at the point (0, -2) the y-intercept. Then from that point run over 4 units and up 3 units to arrive at the point (4, 1). Connect the dots.

OpenStudy (anonymous):

huh?? so the y intercept is (0,-2) and x-intercept os (4,1)?

OpenStudy (saifoo.khan):

Incorrect.

OpenStudy (saifoo.khan):

x is (8/3,0)